CPK messages are initially sent to the CPK mailing list, you can (un)subscribe via this link. You can also follow the service interruption messages via RSS using the link in the title under the RSS icon. If the CPK takes more time to resolve, any updates are published on this website.

For RU wide service interruption see meldingen.ru.nl.

 

Service Interruptions


1415: Clusternode maintenance day - February 6th 2026

Every half year we do clusternode maintenance, with at least a package ugprade and a reboot, but sometimes other maintenance can happen, such as changes in filesystems or network configurations. The upcoming date for this maintenance is February 6th, 2026 (Friday)

Resolved Reports


1422: Self‑creating login broken on DIY

After an upgrade on Feb 3, the DIY service kept an old process running, so self‑service account creation failed. 12 users hit the “account creation failed” error between 12:01 AM and 9:33 AM. The process was restarted, the new version took over and login creation works again. Sorry for the hassle!

1421: Mail Lost between 18:00 2 Feb 2026 and 9:00 3 Feb 2026

We are working on replacing our incoming mailservers and something went wrong. Due to an unexpected block in the mail flow, around 2800 mails which were queued from around 18:00 on Feb 2nd, were lost. The host queuing the mail, was replaced on Feb 3rd, around 8:30 and decommissioned at around 9:00. Unfortunately, the decommissioning also destroyed the disk of the server with the mail queue of around 2800 e-mails. At 18:00 on Feb 2nd, two servers were accepting incoming mail, but only one of the servers was queuing, the other one was running as expected. ...

1420: Certificate expired on Elabftw

Elabftw had an expired certificate, because after the move to a new server, this process was left unmonitored. Both the certificate and the monitoring are now fixed.

1419: Daily backups on peck failed

A combination of small failures in configuration and monitoring resulted in Daily backups not being run on peck between September 1st and last Friday. Backups are now running again, but other than the Monthly backups, there are no intermediate Daily backups for the volumes on server peck. Updates in monitoring/testing are in progress to prevent this kind of oversight in the future.

1418: Elabftw password reset problem

After moving the elabftw site to a new server, it was impossible to send e-mail from the server, like when a user wants to reset their password. The changes that were necessary are in between the host and the docker container running elabftw, while figuring out what was needed also caused some general issues reaching the elabftw webserver.

1417: Network changes websites

To make the network less complicated and more easily manageable, we need to make vlans for the subnets 131.174.8.0/24 and 131.174.9.0/24, this will involve some servers getting a new ip address. This affects firewall rules as well. We must coordinate this with ILS networking and the changes will cause disruptions between a few minutes up to a quarter of an hour. List of sites: alice.science.ru.nl alicecontrol.science.ru.nl coursetimetable.science.ru.nl dmarcreport.science.ru.nl edutools.science.ru.nl eduview.science.ru.nl ...

1416: Radius certificate expired (Eduroam login)

The certificate of the radius server has been replaced this morning, users who use their science account for authenticating to the Eduroam network will not have been able to connect until the certificate was fixed. You may need to ’trust-on-first-use’ or other setting in your device to properly reconnect and trust the certificates, if that still fails.

1414: ldap unavailable

A brief service interruption occurred this morning when one of our LDAP servers refused to start. This affected services that depend on LDAP for user authentication. Services such as Eduroam (using science login for authentication) and GitLab experienced authentication failures during this time. We are investigating the source of the failure to prevent recurrence.

1411: Homeserver upgrade memory

Recently we have experienced some issues with nfs filesystems one of the main suspects was the ZFS filesystem having a lack of cache memory on the machine. In order to improve the situation we will install more memory in the servers. This will cause a downtime of about 10-15 minutes per server. We plan to do this before 10:00 or after 16:00 on Monday Oct 27, 2025. Update 27 Oct 2025 The memory upgrade was concluded around 09:15 on Monday 27 Oct. ...